OpenStudy (anonymous):

Find the GCF of the terms of the polynomial 42 x^5 + 60 x^4

OpenStudy (tkhunny):

Really big hint: 42 = 6*7 60 = 6*10

OpenStudy (anonymous):

6x^6 ?